OpenStudy (anonymous):

I gotta a nice info., its really cool . do U know that while you're sitting still, U're actually traveling around the axis of the earthfaster than a jet plane?????? . . can anybody tell me hows that ? kinda puzzle hah ^_^

OpenStudy (anonymous):

physiciast@ uh, thanks? uh, your a meanie? Uh, thanks for becoming my fan?

OpenStudy (anonymous):

look on google

OpenStudy (anonymous):

I don't always quote Marcel the shell but "Compared to what?" motion is all relative to other objects. The person on the Earth's surface is moving with the earth and of course you get into v=2πr/t whereas a traveler on a plane could be just sitting there not moving at all relative to the plane.

OpenStudy (anonymous):

Compared to the centre of the universe (a theoretical point), you can be screaming across the cosmos at 490,000 mph

OpenStudy (anonymous):

true, it's my fault . I didn't specify for which plane? thats a great job man . okay lets assume that we're talking relative to the person on the earth

OpenStudy (underhill):

The diameter of the earth is 12,756.2 kilometers. It rotates once every 24 hours. Therefore the speed of a person on the equator due to the earth's rotation is\[(diameter \times \pi)\div time\] which equals \[(12,756.2 km \times \pi)\div 24 hours\] which turns out to be\[14.76m/s\]This is the speed due to the earth's rotation; of course, factoring in the earth's revolution around the sun may well certainly bring that speed above that of a jet.

OpenStudy (anonymous):

congrats Underhill you got the point but are you sure about the calculations?

OpenStudy (underhill):

Arghhhhh! My calculator hijacked my solution! ;-) Not. Thanks Physicist Ali - not sure how I got that answer. The real speed of a person on the equator due to the earth's rotation is\[463.83m/s\]